Database last updated on Wed, 04 Feb 2026 09:10:51 +0000 by localhost
Instrument constant changes for the week before 2023250 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 19 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
dt
ed
fm
hg
rs
sl
uv
zs
250
RESOLUTE
#031
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.169)
1
2
16
15
1
7
84
87
(0 sec)
Dead time (1 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B25023.031
1
#031
35
34